OpenStudy (anonymous):

1. Rick used 8 paperclips to put together 4 information packets. What is the greatest number of packets Rick can put together using 32 paperclips?

OpenStudy (anonymous):

??

OpenStudy (triciaal):

you have 32 paperclips for each group of 8 you will have 4 info packets

OpenStudy (susanna):

make a proportion 8/4=32/x cross multiply 8x=128 128/8=x x=16